OpenStudy (anonymous):

Dan and Frank are roommates. They both leave their house on their bicycles at the same time. Dan rides west for 4 miles and then south for 3 miles. Frank rides east for 1 mile and then north for 2 miles. How far apart are they when they both arrive at their destinations?

OpenStudy (anonymous):

can you draw the situation? use the "draw" button below to make a sketch of the problem....

OpenStudy (anonymous):

According to the question below, how far apart are Dan and Frank? Round your answer to two decimal places.

OpenStudy (anonymous):



OpenStudy (anonymous):

3^2+4^2=c^2 c=5 q^2=1^2+2^2 q=~2.24 2.24+5=7.24 6.24 miles apart

OpenStudy (anonymous):

sorry, 7.24

OpenStudy (anonymous):

you got it?